If you travel 798 miles at an average speed of 95 miles per hour, what speed do you have to be going to average 95 miles per hour. Assuming you only stopped 3 times for 10 minutes each time.
Answer by lwsshak3(11628)

If you travel 798 miles at an average speed of 95 miles per hour, what speed do you have to be going to average 95 miles per hour. Assuming you only stopped 3 times for 10 minutes each time.
***
let x=travel time on the road
x+.5 hr=total travel time including 3-10 min stops
distance/travel time=speed
798/x+.5=95
798=95x+47.5
x=7.9 hrs
speed on the road to average 95 mph=798/7.9=101 mph